OpenStudy (anonymous):

Math question help @austin1617

OpenStudy (anonymous):

Where's it at

OpenStudy (anonymous):

it wont let me post the picture in the actual question

OpenStudy (anonymous):

Maximum profit: $70 Trick Skis:6 Slalom Skis:10

OpenStudy (anonymous):

5*6=30 10*4=40 40+30=70

OpenStudy (anonymous):

ok i get it thanks

OpenStudy (anonymous):

Need anymore help?

OpenStudy (anonymous):

ill take as much help as i can get the practice test has 9 question and honestly dont understand any of them

OpenStudy (anonymous):

actually for this one i still dont get how you got those numbers

OpenStudy (anonymous):

Where that question at.

OpenStudy (anonymous):

30/5=6 40/4=10

OpenStudy (anonymous):

sorry but for this question wouldnt it be 1400,0, and 35?

OpenStudy (anonymous):

Then I dont know much about the question then if your getting different answers.

OpenStudy (anonymous):

ok ill try and figure this one out on my own can you help me with a diffrent one

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I don't know if there's a certain equation.

OpenStudy (anonymous):

i dont know either

OpenStudy (anonymous):

You must of learned it somewhere.

OpenStudy (anonymous):

i haven't my school doesn't have much help other then tests

OpenStudy (anonymous):

Well that's a dumb way for tests. You should of learned from an assignment or quiz before the test. They can't give you tests when you don't even know how to do them...

OpenStudy (anonymous):

ya i know my school sucks

OpenStudy (anonymous):

Don't you have anymore questions you want me to see to help with?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well tell that to your school lol.

OpenStudy (anonymous):

they dont listen

OpenStudy (anonymous):

That your educational system with Tests are whack

OpenStudy (anonymous):

ive tried to even the principle sucks in my school

OpenStudy (anonymous):

r+s-4t=-1 4s+t=-3 t=1 4+(-1)-4(1)=-1 4(-1)+1=-3 t=-1 r=4 s=-1 t=1 @rika3307

OpenStudy (anonymous):

@rika3307

OpenStudy (anonymous):

awesome thank u so much

OpenStudy (anonymous):

Anymore more help? @rika3307

OpenStudy (anonymous):

trying to figure these out b4 a huge unit test was the worst and yes i do have a few more if im able to get the help

OpenStudy (anonymous):

This question is hard too.

OpenStudy (anonymous):

could you help me out with a different one its the last one

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

x+5y=26 4x-y=20 6+5(4)=26 4(6)-4=20 x=6 y=4

OpenStudy (anonymous):